# Safe Learning for Contact-Rich Robot Tasks: A Survey from Classical Learning-Based Methods to Safe Foundation Models

### v1

@@ 3.3.5 Tactile Sensing
There exist many types of tactile sensors. In robot learning, the most widely used are vision-based tactile sensors such as the GelSight family [yuan2017gelsight, wang2021gelsight], as well as low-cost, easily arrayable tactile sensor designs such as flexible tactile sensor array [huang20243d]. There has also been progress in vision-based tactile sensors with task-adaptive geometries, such as hemispherical and finger-shaped designs [11024242, sun2022soft, 9811966]. High-resolution vision-based tactile sensing provides high-resolution, local information about contact that complements global vision and low-dimensional F/T measurements. Rather than only measuring a single wrench at the wrist, tactile skins and fingertip arrays reveal pressure distributions, shear and incipient slip at the exact points of interaction, making it possible to detect unsafe contact patterns, such as concentrated loads on corners or sliding on fragile surfaces, before they escalate into damage or failure [31, 43, 71]. For safe learning, these rich contact cues are typically fed directly into the policy alongside proprioceptive state, so that both exploration and execution can be guided away from behaviours that produce sharp spikes or unstable contact transitions [71, 106].
@@
One group of methods uses tactile arrays as the main channel for encoding task success and safe contact configurations. In furniture assembly, for example, tactile ensembles distributed across fingers or grippers are used to characterise how surfaces, edges and fasteners behaviors during successful insertions or alignments. Policies are then trained to reproduce these tactile signatures while avoiding contact patterns associated with jamming or over-tightening [43]. Gentle manipulation frameworks similarly treat touch as the primary feedback for regulating interaction intensity. Policies observe proprioception and tactile events, and are rewarded for acquiring task-relevant information while keeping per-taxel pressures and impact magnitudes low, which naturally steers exploration toward soft, sliding contacts rather than hard impacts [106]. In both cases, safety is not an external constraint but an emergent property of how tactile patterns are encoded in the observation and reward.
Tactile sensing also plays a key role in variable-impedance control for contact-rich motion along surfaces. In aerial or lightweight manipulators that slide along uneven terrain or structural elements, the policy often receives only proprioceptive state and tactile readings, without full F/T information at the base [71]. The tactile stream indicates whether contact is lost, sticking, or concentrated on an unsafe subset of the surface, while the proprioceptive state captures global motion. Variable-impedance updates are then conditioned on these signals, increasing stiffness to maintain stable contact when support is good, and softening or backing off when tactile cues reveal high local pressure or impending slip [71]. This coupling allows robots with limited payload and sensing to exploit contact safely, even when the exact environment model is unknown.
@@
A further set of approaches embeds tactile input within multimodal visual–haptic–proprioceptive representations. Hierarchical variable-impedance controllers, for instance, build high-level policies on visual features that capture object pose and scene layout, while lower-level modules refine actions using haptic channels that may include both force/torque and tactile measurements [31]. In such designs, vision plans where and when contact should occur, while tactile feedback confirms whether contact is distributed and stable in the way the planner expects. More recent multimodal RL systems explicitly combine visual observations, tactile signals and joint state into a shared latent space, and train adaptive policies that allocate different roles to each modality. Vision is used to anticipate geometric constraints, tactile cues to detect unsafe local pressure or slip, and proprioception to enforce kinodynamic limits [86]. By tying safety-related decisions to tactile cues rather than to coarse force thresholds alone, these methods achieve more nuanced and robust protection for delicate objects, complex assemblies and human collaborators.
@@ 3.4 Data Acquisition
Data matters for policy learning, especially for a safe policy where negative samples play a significant role in training the agent to be aware of what unsafe scenarios are. However, acquiring such data is time-consuming and labor-intensive. Crucially, unlike the vast text and image corpora scraped from the Internet to train LLMs and VLAs, physical interaction data, specifically high-fidelity contact forces, torques, and safety-critical failure modes, cannot be obtained from the web. This scarcity of contact data in foundation models creates a significant gap in their ability to reason about physical dynamics and safety. Consequently, physical data collection remains inevitable and irreplaceable for contact-rich tasks, despite efforts to leverage simulation and internet-scale proxies. As shown in the data pyramid Fig.4, real-world contact data sits at the apex—scarce but essential. And in Fig.4 we retain only the most representative concepts in each layer to highlight the progression rather than exhaustively listing all sources. Within each layer, the listed keywords are ordered by increasing contact fidelity and safety relevance: for real-world data, from generic task execution logs to explicit safety and failure cases; for simulation, from generic randomized environments to simulation elements that most directly support sim-to-real transfer; and for internet and shared datasets, from general web-scale corpora to task-specific real-robot benchmarks.
@@ 3.4.1 Simulation-Based Data Generation
Simulated environments provide a safe, scalable, and cost-effective way to generate large volumes of data for contact-rich tasks without risking damage to hardware or fixtures. Physics engines such as MuJoCo [todorov2012mujoco], MuJoCo Playground [zakka2025mujoco], PyBullet [coumans2016pybullet], and Isaac Gym [makoviychuk2021isaac] enable learning-based agents to interact with objects under diverse contact conditions, often with ground-truth access to forces, velocities, object poses, and other latent quantities that are difficult or expensive to measure on real systems. These simulators support large-scale parallel rollout, accelerating the training of data-hungry methods such as deep RL and imitation learning for contact-rich manipulation.
@@ 3.4.2 Real-World Data Collection
Real-world data collection remains a cornerstone of safe learning, as it captures the complexities and uncertainties of physical interactions that simulations may not fully replicate. This involves deploying robots in controlled environments or real-world settings to gather data on contact forces, object dynamics, and task performance under various conditions. However, collecting high-quality data in the real world is often time-consuming, expensive, and potentially unsafe, especially in contact-rich tasks where excessive forces or unexpected contacts can damage hardware or pose safety risks to humans. To address these challenges, researchers often employ techniques such as teleoperation, where human operators guide the robot through tasks while collecting data on safe and unsafe interactions [mower2023ros, 354]. Additionally, safety constraints can be integrated into the data collection process, ensuring that the robot operates within predefined limits to avoid damaging itself or its environment [5, 18]. Furthermore, real-world data can be augmented with simulation data through techniques like domain adaptation or transfer learning, allowing policies trained in simulation to be fine-tuned on real-world data [zhang2024srl, berrocal2024evaluating].
@@
MuJoCo Manipulation, Locomotion [zhang2024srl]
MuJoCo XLA/MJX Manipulation, Locomotion [zhang2024srl]
Genesis Soft-body, Articulated [Genesis]
PyBullet Manipulation, Grasping [mower2023ros, 354]
Isaac Gym RL Training, Manipulation [makoviychuk2021isaac]
Isaac Sim Photorealistic, Digital Twin [noauthor_isaac_nodate, berrocal2024evaluating]
@@
Orbit Interactive Robot Learning [mittal2023orbit]
Drake Model-based Control [tedrake2014drake]
Factory RL, Assembly [narang2022factory]
@@
RoboVerse Manipulation, RL Benchmarks [geng2025roboverse]
RoboCasa Household Tasks, Manipulation [nasiriany2024robocasa]
RoboSuite Manipulation, RL Benchmarks [zhu2020robosuite]
ManiSkill Manipulation, RL Benchmarks [mu2maniskill, gumaniskill2]
Meta-World Manipulation, RL Benchmarks [yu2020meta]
OpenAI Gym RL Benchmarks, Manipulation [brockman2016openai]
RLBench Manipulation, RL Benchmarks [james2020rlbench]
Stable Baselines RL Benchmarks, Manipulation [raffin2021stable]
Gymnasium RL Benchmarks, Manipulation [towers2024gymnasium]
@@
Robust gymnasium Safety in RL, Robotics [gu2025robust]
Safe-Control-Gym Safety in RL, Control [yuan2022safe]
@@ 3.4.3 Hybrid and Self-Supervised Data Acquisition
Hybrid approaches combine simulation and real-world data to leverage the strengths of both modalities. By pre-training policies in simulation and then fine-tuning them on real-world data, researchers can achieve better generalization and robustness in contact-rich tasks. This approach allows for the efficient exploration of a wide range of scenarios in simulation while still adapting to the nuances of real-world interactions [berrocal2024evaluating]. Self-supervised learning techniques can also be employed to generate additional training signals from the robot’s own interactions, allowing it to learn from its experiences without requiring extensive labeled data [zhang2024srl]. For example, robots can learn to predict future states or contact forces based on their own sensory inputs, creating a form of intrinsic motivation that drives exploration and learning in contact-rich environments [zhang2024srl]. These hybrid and self-supervised approaches can help bridge the reality gap and improve the robustness of learned policies in contact-rich tasks, enabling them to generalize better to real-world scenarios. By combining simulated and real-world data, researchers can create more comprehensive training datasets that capture a wider range of contact scenarios, leading to safer and more effective policies for robotic manipulation. For instance, researchers may pre-train policies in simulation and then fine-tune them on real-world data, or use simulated data to augment limited real-world datasets [berrocal2024evaluating]. Self-supervised learning techniques can also be employed to generate additional training signals from the robot’s own interactions, allowing it to learn from its experiences without requiring extensive labeled data [zhang2024srl]. These approaches can help bridge the reality gap and improve the robustness of learned policies in contact-rich tasks, enabling them to generalize better to real-world scenarios.
@@ 3.5 Simulation Environment and Benchmarks
Simulation plays a critical role in the development and evaluation of safe learning algorithms for contact-rich robotic tasks. These tasks often involve complex dynamics, high-dimensional state spaces, and the potential for hardware damage, making direct deployment on real robots risky and costly during early stages of learning. Simulated environments provide a safe, controllable, and repeatable platform to test algorithms, enabling rapid iteration and the generation of large-scale data without physical wear.  [long2025survey]
@@ 3.5.1 Simulation Environment
We group commonly used tools into a richer taxonomy reflecting the physics core, parallelism, sensing realism, and intended evaluation protocols. We highlight the perspective of contact-rich tasks in this survey, readers can find more comparative details in [berrocal2024evaluating].
@@
Rigid-body dynamics engines. CPU engines such as DART (via robot_dart) [251] emphasize accurate contact, joint limits, and articulated constraints. PyBullet [93, 139, 154, 254, 263, 302, 185] remains a widely used open-source engine with strong community support. MuJoCo, one of the most popular simulators for contact-rich tasks, is known for its inherent contact modeling and fast simulation speed. It powers standard benchmarks such as RoboSuite [201] and Safety Gym [262], and is extensively used in contact-rich manipulation tasks including insertion [85, 104, 279], assembly [99, 174], and surface interaction [66, 151]. Furthermore, it supports research in safe exploration and control [29, 65, 91, 129], as well as data collection and skill learning [18, 37, 74, 77, 178]. It is also widely adopted in non-prehensile and articulated object manipulation [7, 81, 94, 145], dynamic and locomotion tasks [100, 101], and complex media interaction [92, 106]. Additional applications span diverse contact-rich scenarios [108, 109, 119, 120, 125, 126, 128, 133, 135, 150, 159, 170, 173, 190, 197, 289].
GPU engines leverage massive parallelism to accelerate data collection for RL. The Isaac ecosystem is widely adopted, including Isaac Gym and Isaac Sim, along with their modular frameworks like Factory [narang2022factory], Orbit [mittal2023orbit], Isaac Lab [372], and RoboVerse [geng2025roboverse]. These platforms support large-scale training for contact-rich tasks [10, 22, 207, 240, 252, 298]. Recently, MuJoCo MJX [zakka2025mujoco] has also introduced GPU-accelerated physics to the MuJoCo ecosystem.
@@ 3.5.2 Benchmarks
General manipulation suites. RoboSuite [201] standardizes robot models and canonical tasks (insertion, door opening), serving as a primary testbed for modular design. Meta-World [yu2020meta] and RLBench [james2020rlbench] offer multi-task breadth, though often with simplified contact physics. Recently, RoboVerse [geng2025roboverse] and RoboCasa [nasiriany2024robocasa] have scaled these to large-scale, scene-level manipulation tasks using foundation models.
Contact-intensive and generalizable benchmarks. To address the reality gap in contact physics, ManiSkill [mu2maniskill] and ManiSkill2 [gumaniskill2] feature diverse articulated objects and high-fidelity physics for learning generalizable policies. Factory [narang2022factory] provides physically accurate assembly tasks (nut-bolt threading) specifically designed for testing contact-rich policies in Isaac Gym.
Safety-oriented benchmarks. Safety Gym and its successor Safety Gymnasium [262, 152] are the de facto standards for safe RL, defining constrained navigation and manipulation tasks with hazard maps and cost signals. Beyond penalty-based safety, Safe-Control-Gym [yuan2022safe] integrates control-theoretic constraints (symbolic priors, partial observability) for evaluating certified safety methods. Robust Gymnasium [gu2025robust] specifically targets safety under distributional shifts and adversarial disturbances, critical for robust sim-to-real transfer.

### v2

@@ 3.3.5 Tactile Sensing
There exist many types of tactile sensors, which provide direct measurements of contact states (e.g., contact location, normal/shear forces, and incipient slip) and thus play a critical role in safe execution of contact-rich task. In robot learning, the most widely used are vision-based tactile sensors such as the GelSight family [361, 321], as well as low-cost, easily arrayable tactile sensor designs such as flexible tactile sensor array [111]. There has also been progress in vision-based tactile sensors with task-adaptive geometries, such as hemispherical and finger-shaped designs [125, 288, 67]. High-resolution vision-based tactile sensing provides high-resolution, local information about contact that complements global vision and low-dimensional F/T measurements. Rather than only measuring a single wrench at the wrist, tactile skins and fingertip arrays reveal pressure distributions, shear and incipient slip at the exact points of interaction, making it possible to detect unsafe contact patterns, such as concentrated loads on corners or sliding on fragile surfaces, before they escalate into damage or failure [164, 172, 379]. Beyond conventional parallel-jaw grippers and tactile-equipped dexterous hands [390, 27, 292], other end-effectors—such as multi-DoF (e.g.,5-DoF) grippers—can enable manipulation primitives like cable/rope disentangling and thin-card flipping [394]. These designs broaden the repertoire of feasible manipulation tasks and could be promising to combine with safety modules or learned control policies in future work. For safe learning, these rich contact cues are typically fed directly into the policy alongside proprioceptive state, so that both exploration and execution can be guided away from behaviours that produce sharp spikes or unstable contact transitions [379, 114].
@@
One group of methods uses tactile arrays as the main channel for encoding task success and safe contact configurations. [401] uses a large-scale visuo-tactile dataset consisting of over 2.6 million vision–touch pairs, collected across multiple environments and covering 43 manipulation tasks. The dataset is used to pretrain a visuo-tactile representation (encoder) as an upstream step, which can later serve as a foundation for downstream work such as safe contact-rich interaction or manipulation. In furniture assembly, for example, tactile ensembles distributed across fingers or grippers are used to characterise how surfaces, edges and fasteners behaviors during successful insertions or alignments. Policies are then trained to reproduce these tactile signatures while avoiding contact patterns associated with jamming or over-tightening [172]. Gentle manipulation frameworks similarly treat touch as the primary feedback for regulating interaction intensity. Policies observe proprioception and tactile events, and are rewarded for acquiring task-relevant information while keeping per-taxel pressures and impact magnitudes low, which naturally steers exploration toward soft, sliding contacts rather than hard impacts [114]. In both cases, safety is not an external constraint but an emergent property of how tactile patterns are encoded in the observation and reward.
Tactile sensing plays a key role in variable-impedance control for contact-rich motion along surfaces. In aerial or lightweight manipulators that slide along uneven terrain or structural elements, the policy often receives only proprioceptive state and tactile readings, without full base F/T sensing [379]. The tactile stream reveals contact loss, stick–slip, and localized high-pressure/unsafe contact, while proprioception captures global motion; impedance can then be updated online—stiffening when contact is well supported and softening/backing off when tactile cues indicate high pressure or impending slip [379]. Beyond aerial sliding, tactile/visuo-tactile interaction controllers explicitly use curvature/alignment monitoring to adapt stiffness and force regulation when exploring unknown 3D curved surfaces, including contact-loss recovery and passivity-aware safety mechanisms [129, 130]. For continuous surface tasks (e.g., cleaning/inspection), closed-loop tactile coverage methods encode the target on point clouds and use task-space impedance with force regulation to maintain contact while covering curved surfaces [28].
@@
A further set of approaches embeds tactile input within multimodal visual haptic proprioceptive representations [376]. Hierarchical variable-impedance controllers, for instance, build high-level policies on visual features that capture object pose and scene layout, while lower-level modules refine actions using haptic channels that may include both force/torque and tactile measurements [164]. In such designs, vision plans where and when contact should occur, while tactile feedback confirms whether contact is distributed and stable in the way the planner expects. More recent multimodal RL systems explicitly combine visual observations, tactile signals and joint state into a shared latent space, and train adaptive policies that allocate different roles to each modality. Vision is used to anticipate geometric constraints, tactile cues to detect unsafe local pressure or slip, and proprioception to enforce kinodynamic limits [311]. By tying safety-related decisions to tactile cues rather than to coarse force thresholds alone, these methods achieve more nuanced and robust protection for delicate objects, complex assemblies and human collaborators.
@@ 3.4 Data Acquisition
Data matters for policy learning, especially for a safe policy where negative samples play a significant role in training the agent to be aware of what unsafe scenarios are. However, acquiring such data is time-consuming and labor-intensive. Crucially, unlike the vast text and image corpora scraped from the Internet to train LLMs and VLAs, physical interaction data, specifically high-fidelity contact forces, torques, and safety-critical failure modes, cannot be obtained from the web. This scarcity of contact data in foundation models creates a significant gap in their ability to reason about physical dynamics and safety. Consequently, physical data collection remains inevitable and irreplaceable for contact-rich tasks, despite efforts to leverage simulation and internet-scale proxies. As shown in the data pyramid in Fig. 6, real-world contact data sits at the apex and is both scarce and essential. And in Fig.6 we retain only the most representative concepts in each layer to highlight the progression rather than exhaustively listing all sources. Within each layer, the listed keywords are ordered by increasing contact fidelity and safety relevance: for real-world data, from generic task execution logs to explicit safety and failure cases; for simulation, from generic randomized environments to simulation elements that most directly support sim-to-real transfer; and for internet and shared datasets, from general web-scale corpora to task-specific real-robot benchmarks.
@@ 3.4.1 Simulation-Based Data Generation
Simulated environments provide a safe, scalable, and cost-effective way to generate large volumes of data for contact-rich tasks without risking damage to hardware or fixtures. Physics engines such as MuJoCo [305], MuJoCo Playground [363], PyBullet [55], Isaac Gym [202] and Isaac Lab [218] enable learning-based agents to interact with objects under diverse contact conditions, often with ground-truth access to forces, velocities, object poses, and other latent quantities that are difficult or expensive to measure on real systems. These simulators support large-scale parallel rollout, accelerating the training of data-hungry methods such as deep RL and imitation learning for contact-rich manipulation.
@@ 3.4.2 Real-World Data Collection
Real-world data collection remains a cornerstone of safe learning, as it captures the complexities and uncertainties of physical interactions that simulations may not fully replicate. This involves deploying robots in controlled environments or real-world settings to gather data on contact forces, object dynamics, and task performance under various conditions. However, collecting high-quality data in the real world is often time-consuming, expensive, and potentially unsafe, especially in contact-rich tasks where excessive forces or unexpected contacts can damage hardware or pose safety risks to humans. To address these challenges, researchers often employ techniques such as teleoperation, where human operators guide the robot through tasks while collecting data on safe and unsafe interactions [220, 192]. Additionally, safety constraints can be integrated into the data collection process, ensuring that the robot operates within predefined limits to avoid damaging itself or its environment [345, 181]. Furthermore, real-world data can be augmented with simulation data through techniques like domain adaptation or transfer learning, allowing policies trained in simulation to be fine-tuned on real-world data [374, 25].
@@
MuJoCo Manipulation, Locomotion [374]
MuJoCo XLA/MJX Manipulation, Locomotion [374]
Genesis Soft-body, Articulated [13]
PyBullet Manipulation, Grasping [220, 192]
Isaac Gym RL Training, Manipulation [202]
Isaac Sim Photorealistic, Digital Twin [227, 25]
@@
Orbit Interactive Robot Learning [219]
Factory RL, Assembly [223]
RoboAssembly ✗? furniture assembly [357]
Drake Model-based Control [299]
@@
RoboVerse Manipulation, RL Benchmarks [82]
RoboCasa Household Tasks, Manipulation [224]
RoboSuite Manipulation, RL Benchmarks [402]
ManiSkill Manipulation, RL Benchmarks [221, 87]
Robot Air Hockey Air Hockey game [180]
Meta-World Manipulation, RL Benchmarks [358]
OpenAI Gym RL Benchmarks, Manipulation [38]
RLBench Manipulation, RL Benchmarks [122]
Stable Baselines RL Benchmarks, Manipulation [243]
Gymnasium RL Benchmarks, Manipulation [308]
@@
Robust gymnasium Safety in RL, Robotics [91]
Safe-Control-Gym Safety in RL, Control [362]
@@ 3.4.3 Hybrid and Self-Supervised Data Acquisition
Hybrid approaches combine simulation and real-world data to leverage the strengths of both modalities. By pre-training policies in simulation and then fine-tuning them on real-world data, researchers can achieve better generalization and robustness in contact-rich tasks. This approach allows for the efficient exploration of a wide range of scenarios in simulation while still adapting to the nuances of real-world interactions [25]. Self-supervised learning techniques can also be employed to generate additional training signals from the robot’s own interactions, allowing it to learn from its experiences without requiring extensive labeled data [374]. For example, robots can learn to predict future states or contact forces based on their own sensory inputs, creating a form of intrinsic motivation that drives exploration and learning in contact-rich environments [374]. These hybrid and self-supervised approaches can help bridge the reality gap and improve the robustness of learned policies in contact-rich tasks, enabling them to generalize better to real-world scenarios. By combining simulated and real-world data, researchers can create more comprehensive training datasets that capture a wider range of contact scenarios, leading to safer and more effective policies for robotic manipulation. For instance, researchers may pre-train policies in simulation and then fine-tune them on real-world data, or use simulated data to augment limited real-world datasets [25]. Self-supervised learning techniques can also be employed to generate additional training signals from the robot’s own interactions, allowing it to learn from its experiences without requiring extensive labeled data [374]. These approaches can help bridge the reality gap and improve the robustness of learned policies in contact-rich tasks, enabling them to generalize better to real-world scenarios.
@@ 3.5 Simulation Environment and Benchmarks
Simulation plays a critical role in the development and evaluation of safe learning algorithms for contact-rich robotic tasks. These tasks often involve complex dynamics, high-dimensional state spaces, and the potential for hardware damage, making direct deployment on real robots risky and costly during early stages of learning. Simulated environments provide a safe, controllable, and repeatable platform to test algorithms, enabling rapid iteration and the generation of large-scale data without physical wear [190].
@@ 3.5.1 Simulation Environment
We group commonly used tools into a richer taxonomy reflecting the physics core, parallelism, sensing realism, and intended evaluation protocols. We highlight the perspective of contact-rich tasks in this survey, readers can find more comparative details in [25].
@@
Rigid-body dynamics engines. CPU engines such as DART (via robot_dart) [209] emphasize accurate contact, joint limits, and articulated constraints. PyBullet [71, 267, 268, 135, 360, 235, 284] remains a widely used open-source engine with strong community support. MuJoCo, which is one of the most popular simulators for contact-rich tasks for its inherent contact modeling and fast simulation speed. It powers standard benchmarks such as RoboSuite [186] and Safety Gym [278], and is extensively used in contact-rich manipulation tasks including insertion [280, 23, 256], assembly [347, 97], and surface interaction [399, 158]. Furthermore, it supports research in safe exploration and control [132, 366, 301, 33], as well as data collection and skill learning [181, 381, 175, 22, 244]. It is also widely adopted in non-prehensile and articulated object manipulation [1, 137, 331, 382], dynamic and locomotion tasks [14, 275], and complex media interaction [146, 114]. Additional applications span diverse contact-rich scenarios [213, 53, 43, 250, 42, 395, 72, 142, 337, 136, 152, 153, 380, 147, 12, 300].
GPU engines leverage massive parallelism to accelerate data collection for RL. The Isaac ecosystem is widely adopted, including Isaac Gym and Isaac Sim, along with their modular frameworks like Factory [223], Orbit [219], Isaac Lab [228], and RoboVerse [82]. These platforms support large-scale training for contact-rich tasks [226, 138, 100, 295, 69, 51]. Recently, MuJoCo MJX [363] has also introduced GPU-accelerated physics to the MuJoCo ecosystem.
@@ 3.5.2 Benchmarks
General manipulation suites. RoboSuite [186] standardizes robot models and canonical tasks (insertion, door opening), serving as a primary testbed for modular design. Meta-World [358] and RLBench [122] offer multi-task breadth, though often with simplified contact physics. Recently, RoboVerse [82] and RoboCasa [224] have scaled these to large-scale, scene-level manipulation tasks using foundation models.
Contact-intensive and generalizable benchmarks. To address the reality gap in contact physics, ManiSkill [221] and ManiSkill2 [87] feature diverse articulated objects and high-fidelity physics for learning generalizable policies, including contact-rich tasks (plug charger, peg insertion, etc.). Factory [223] provides physically accurate assembly tasks (nut-bolt threading) specifically designed for testing contact-rich policies in Isaac Gym.
Safety-oriented benchmarks. Safety Gym and its successor Safety Gymnasium [278, 123] are the de facto standards for safe RL, defining constrained navigation and manipulation tasks with hazard maps and cost signals. Beyond penalty-based safety, Safe-Control-Gym [362] integrates control-theoretic constraints (symbolic priors, partial observability) for evaluating certified safety methods. Robust Gymnasium [91] specifically targets safety under distributional shifts and adversarial disturbances, critical for robust sim-to-real transfer.